OpenStudy (anonymous):

You must show all your work on each of the following questions. 1. Solve 5x^2 = –45x. 2. Solve x^2 = 14x – 49. 3. Solve x^2 + 18 = –9x. 4. Solve 4x^2 = 12x + 40.

OpenStudy (anonymous):

\[5x^2 = –45x\] \[5 x^2+45 x = 0\] \[5x(x + 9) = 0\] \[5x = 0\] \[x = 0\] \[(x + 9) = 0\] \[x = -9\]

OpenStudy (anonymous):

For all of these problems, you can rearrange the equations in the form \[ax^2 + bx + c = 0\] From there, all you need to do is factorise. The first one, for example, rearranges to 5x^2+45x = 0, which factorises to 5x(x+9)=0. You can see that either 5x=0, or x+9=0, from which you can get two solutions. If you can't factorise the equation, use the quadratic formula.

OpenStudy (anonymous):

\[x^2 = 14x – 49\] \[x^2-14 x+49 = 0\] \[(x - 7) (x - 7) = 0\] \[(x - 7) = 0\] \[x = 7\] \[(x - 7) = 0\] \[x = 7\] so x = 7

OpenStudy (anonymous):

\[x^2 + 18 = –9x\] \[x^2+9 x+18 = 0\] \[(x+3) (x+6) = 0\] \[x = -6\] \[x = -3\]

OpenStudy (anonymous):

do you get the hang of it now? @touseii45

OpenStudy (anonymous):

@some_someone A bit

OpenStudy (anonymous):

hmm well what @Fruitbasket wrote is a good explanation, might want to read it :)

OpenStudy (anonymous):

@some_someone why does the answer become the opposite of what came out? like in the last one you put (x+3) (x+6) = 0 but got -3 and -6 rather than 3 and 6. Why is that?

OpenStudy (anonymous):



OpenStudy (anonymous):

ohh alright. now that makes sense. i tried doing the last one but i cant seem to factor it right. Like i cant find the right ones for it. @some_someone

OpenStudy (anonymous):

\[4x^2 = 12x + 40\] \[4 x^2-12 x-40 = 0\] use the quadratic formula.

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

and in your case it would be like this: \[ax^2 + bx + c = 0\] \[4x^2 - 12x - 40 = 0\] where: a = 4 b = -12 c = -40

OpenStudy (anonymous):

\[x = \frac{ -12 \pm \sqrt{-12^2 - 4(4)(-40)} }{ 2(4) } \]

OpenStudy (anonymous):

so its something like this?? @some_someone

OpenStudy (anonymous):

\[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] \[x = \frac{ -(-12) \pm \sqrt{(-12)^2 - 4(4)(-40)} }{ 2(4) }\]

OpenStudy (anonymous):

your is close but not correct @touseii45

OpenStudy (anonymous):

so after all the math x = 5 right? @some_someone

OpenStudy (anonymous):

that is only one solution, the equation has to solutions.

OpenStudy (anonymous):

did you seperate them into two, since its \[\pm \]

OpenStudy (anonymous):

so x = 5 and x = -2 ? @some_someone

OpenStudy (anonymous):

precisely :) good job :D

OpenStudy (anonymous):

Thank you so much!! @some_someone

OpenStudy (anonymous):

yw :)